# Denis Roio Dyne.org &Rebecca Selvaggini

### v2

@@ Abstract
It is of critical importance to design digital identity systems that ensure the privacy of citizens as well as protecting them from issuer corruption. We aim to solve this issue and propose a method for selective disclosure and privacy preserving revocation of digital credentials, using the unique homomorphic characteristics of second order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures. Our approach ensures that users can selectively reveal credentials signed by a certain issuer, which can be interactively revoked by a quorum of other agreeing issuers without revealing the identity of users. Our goal is to protect users from issuer corruption by requiring collective agreement among multiple revocation issuers.
@@ 1.2 Threats considered
We believe that anonymous revocations are a condicio sine qua non to guarantee a sufficient level of privacy in digital identities and credentials. There is no privacy-preserving revocation system designed, either in EUDI-ARF, or in W3C-VC and BBS+. In case the choice of strategy for revocation is left open to developers, the risk for major privacy breaches may occur, for example with the adoption of public status lists [7]. The hypotetical use of a certificate status lists (CRL) presents issues related primarily to privacy [8], because sensitive information about holders leaks from the list. This problem is partially mitigated by expiration dates, in cases where credentials can be short-lived (typically less important credentials), but not applicable with digital identification documents such as ID, driving license, passport and social security numbers, which typically have longer or no expiration time. Future plans for the national standards we are observing include the adoption of “Bitstring” status lists [9] which may grant a degree of privacy. Our approach is that of designing a privacy-preserving revocation mechanism to remove the leak of holder’s information and allow to delegate the governance of revocations to multiple revocation issuers which may be different from the credential issuer.
@@
We briefly round up on feature differences between the named selective disclosure cryptographic schemes, as shown in table 1, mainly distinguishing between three fundamental features:
@@
Where wip is mentioned meaning work in progress on adoption of bitstring status lists for unlinkable revocations.
2 Overview
@@ 2.1 Key contributions
The cryptographic scheme described in this paper, named SD-BLS for brevity, implements some of the properties of the SD-JWT scheme (which can be easily encoded in mDOC format) and proposes a novel cryptographic approach to similar data structures. Furthermore SD-BLS proposes novel anonymous cryptographic revocation flow for verifiable credentials, that aims to solve the privacy issues posed by status and revocation lists.
@@
Similarly to the SD-JWT and mDOC formats, SD-BLS produces an array of claims: the elements of the array are individually signed by the issuer. In SD-BLS the signature(s) replace the HMAC. Using the homomorphic properties of this curve SD-BLS enables the holder to selectively disclose only certain issuer signed claims and produce a presentation that minimizes private information given to verifiers.
@@
SD-BLS proposes a novel approach to credential revocation: the data published by the revocation issuer will produce cryptographic material that contains no information about the credential holders. The cryptographic revocation material allows anyone to verify if an SD-BLS proof produced by a credential holder has been revoked by one or multiple issuers. The unlinkability, and thus anonymity of the cryptographic revocation, allows the revocation issuer to share revocations in public and allows anyone to verify if credentials have been revoked.
@@ 2.2 Applications
The focus of the EUDI-ARF specifications is identity documents: it defines mechanisms and data structures to issue a Personal Identification (PID) as well as digital driving licenses. Similarly, the US government is experimenting with W3C-VC and mDOC for cross-states interoperable driving licenses. SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and anonymous revocation.
@@
  • A verifiable presentation can be used by a smart-contract to verify if one or more holder’s claim match the requirement needed to process a transaction.
@@ 3.1 Notations and assumptions
We also require defining the notion of a cryptographic pairing. That is defined as a function $e:\mathbb{G}_{2}\times\mathbb{G}_{1}\to\mathbb{G}_{T}$, where $\mathbb{G}_{1},\mathbb{G}_{2}$ and $\mathbb{G}_{T}$ are all groups of same order $n$, such that satisfies the following properties:
@@
    Bilinearity, i.e. given $P_{1},Q_{1}\in\mathbb{G}_{1}$ and $P_{2},Q_{2}\in\mathbb{G}_{2}$, we have
@@
    Non-degeneracy, meaning that for all $g_{1}\in\mathbb{G}_{1},g_{2}\in\mathbb{G}_{2}$, $e(g_{2},g_{1})\neq 1_{\mathbb{G}_{T}}$, the identity element of the group $\mathbb{G}_{T}$;
@@
    $\mathbb{G}_{1}\neq\mathbb{G}_{2}$, and moreover, that there exist no efficient homomorphism between $\mathbb{G}_{1}$ and $\mathbb{G}_{2}$.
For the purpose of our protocol we will have $\mathbb{G}_{1}=E$ and $\mathbb{G}_{2}=E_{T}$, and $\mathbb{G}_{T}\subset\mathbb{F}_{p^{12}}$ is the subgroup containing the $n$-th roots of unity, where $n$ is the order of the groups $E$ and $E_{T}$. Instead $e:E_{T}\times E\to\mathbb{G}_{T}$ is the Miller pairing, which in our work is encoded as the method miller(ECP2 P, ECP Q).
@@ 3.2 Issuance
As for other well known algorithms BLS signing will work following three main steps:
  • Key Generation phase. For an issuer who wants to sign a credential $m$, a secret key $sk$ is a random number chosen uniformly in $\mathbb{F}_{n}$, where $n$ is the order of the groups $\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T}$. The corresponding public key $pk$ is the element $sk\cdot G_{2}\in E_{T}$;
  • Signing phase. The credential $m$ is first hashed into the point $U\in E$, which in our scheme is done by the method hashtopoint; the related signature is then given by $\sigma=sk\cdot U$;
  • Verification phase. For an other user that wants to verify the authenticity and the integrity of the message $m$, it needs to
@@
  • Proof of correctness By using the definitions of the elements involved and exploiting the property of the pairing $e$ we have
    $\begin{split}e(pk,U)&=e(sk\cdot G_{2},U)\\ &=e(G_{2},U)^{sk}\\ &=e(G_{2},sk\cdot U)\\ &=e(G_{2},\sigma)\end{split}$
@@
We create a new secret key $rev$ with public key $r$ and signature of the claim $\sigma_{rev}$ to add to the issuer signature:
$\begin{split}r&=rev\cdot G_{2}\\ \sigma_{rev}&=rev\cdot U\\ \sigma&=sk\cdot U+\sigma_{rev}\\ \end{split}$
@@
$\mathcal{C}=\big{\{}\{id,r,\sigma\}:id\in claims\big{\}}$
At the end of this phase the holder is sent the signed claims to be stored in a private wallet, while the issuer stores a list of revocations, associated with an identifier of the holder and the specific credential, into a private database that can be used later to issue revocations.
@@ 3.3 Presentation
When the holder needs to disclose a particular claim $id$ it is sufficient to send the signed claim set $\{id,r,\sigma\}$.
@@ 3.4 Verification
Credential verification is made by checking the presented issuer’s signature. In order to do that the credential issuer’s public key must be added to the revocation public key.
As explained in section 3.3 we can consider the credential proof as a collection of tables of the following form:
$c_{id}=\{id,r,\sigma\}$
where $\sigma$ and $r$ are respectively the signature of the string $id$ and the revocation public key. We can check the validity of the presented claim computing the key
@@
3.4.1 Proof of correctness
The signature $\sigma$ is given by
$\begin{split}\sigma&=sign(A.sk,id)+sign(rev,id)\\ &=A.skU+revU\end{split}$
where $U$ is the mapping of the string $id$ in the group $G_{1}$.
@@ 3.4.1 Proof of correctness
This concludes the first verification phase. If the given presentation is valid, then the verifier should proceed to check the revocation status of the credential as explained below.
@@ 3.5 Revocation
Given an element $c_{id}=\{id,r,\sigma\}$ of the credential presentation and the Issuer public key $A.pk$, if the claim id is present in the revocation list, we can take the corresponding revocation private key $rev$.
We can verify the validity of the revocation by checking if the revocation public key presented by the holder match:
@@ 3.6 Threshold Revocation
Such a revocation dealer will be in contact with credential issuers for the signature of credentials: it will create the $rev$ revocation key while concealing it from them. The trustee then proceeds creating the $\sigma_{rev}=rev\cdot U$ revocation signature and the $r=rev\cdot G_{2}$ public key as detailed in the issuance phase (3.2). Finally the revocation trustee will distribute $\sigma_{rev}$ and $r$ to issuers for further processing into the final signed credential.
The revocation trustee (dealer) will then proceed asynchronously to split the revocation key into shares using a public verifiable secret sharing (PVSS [11]) implementation and distribute these shares to all revocation issuers. In order to issue a revocation, a configurable quorum of peers among the revocation issuers will need to reconstruct the secret revocation key and publish it.
@@ 4 Benchmarks
Figure 1: Speed comparison of issue and verify
Based on our benchmarks, the resulting data objects sizes, excluding the signed material which can be either of variable length or fixed length hashes, are:
  • Proof: 145 Bytes
@@
The computational cost grows for the verification phase with the presence of a cryptographic revocations list. We assume that revocations are grouped by claim type or credential structure, for instance all revocations of a certain common claim (e.g.: "is above 18" or "is a Danish citizen") for any holder are published in the same list: this will not degrade the level of privacy and will group some revocations in smaller lists.
Lab measurements of the time taken by a single proof verification process to operate on a growing number of revocations demonstrate that there is a linear growth, as shown in figure 2.
@@ 5.1 Security considerations
The revocations database is privacy and corruption sensitive (by our previous definition of issuer corruption) and it should be securely stored by each Issuer. This is mitigated by the adoption of threshold revocation. When using threshold revocation there is still the need for a revocation trustee who may be a single point of failure if allowed to publish revocation keys dishonestly collected during the process of credential issuance.
BLS signatures and the proof system obtained with credentials are considered secure by assuming the existence of random oracles [12], together with the decisional Diffie-Hellman Problem (DDH) [13], the external Diffie-Hellman Problem (XDH), and with the Lysyanskaya-Rivest-Sahai-Wol Problem (LRSW) [14], which are connected to the Discrete Logarithm.
@@
The future growth of quantum-computing technologies may be able to overcome the Discrete Logarithmic assumptions by qualitatively different computational means and SD-BLS may then be vulnerable to quantum-computing attacks. However this is speculative reasoning on what we can expect from the future.
The SD-BLS implementation we present in this paper doesn’t integrate any protection against replay-attacks, nor implements timestamps and expiration. We believe security features protecting from replay-attacks can be easily added and adapted from the context of deployment. I absence of session management providing such protections, any real-world use of SD-BLS should contemplate adding a signed timestamp to presentations and expiration checks on verifications.
@@ 5.2 Future development
EUDI-ARF dictates that the holder’s secret key generation and signatures must occur inside a trusted platform module (TPM). For mobile devices, this limits the secret keys and signatures to those offered, via proprietary APIs by the mobile OS, namely RSA (multiple flavours) and ECDSA on the secp256r1 curve. Currently the TPMs APIs supported by Android and iOS do not support BLS 12381 key generation or signature.
@@
The SD-BLS math is fully compatible with ETH2.0 and can be computed inside an Ethereum VM. A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [20]) that want to authenticate peers using the selective disclosure of verifiable credentials instead of a key based proof of possession.

### v4

@@ Abstract
Ensuring privacy and protection from issuer corruption in digital identity systems is crucial. We propose a method for selective disclosure and privacy-preserving revocation of digital credentials using second-order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures. We make holders able to present proofs of possession of selected credentials without disclosing them, and we protect their presentations from replay attacks. Revocations may be distributed among multiple revocation issuers using publicly verifiable secret sharing (PVSS) and activated only by configurable consensus, ensuring robust protection against issuer corruption. Our system’s unique design enables fast revocation checks allowing the use of large revocation lists.
@@ 1.2 Threats considered
We believe that anonymous revocations are a condicio sine qua non to guarantee a sufficient level of privacy in digital identities and credentials. There is no privacy-preserving revocation system designed, either in EUDI-ARF, or in W3C-VC and BBS+. In case the choice of strategy for revocation is left open to developers, the risk for major privacy breaches may occur, for example with the adoption of public status lists [7]. The hypotetical use of a certificate status lists (CRL) presents issues related primarily to privacy [8], because sensitive information about holders leaks from the list. This problem is partially mitigated by expiration dates, in cases where credentials can be short-lived (typically less important credentials), but not applicable with digital identification documents such as ID, driving license, passport and social security numbers, which typically have longer or no expiration time. Future plans for the national standards we are observing include the adoption of “Bitstring” status lists [9] which may grant a degree of privacy. In SD-BLS we design a privacy-preserving revocation mechanism to remove the leak of holder’s information and delegate the governance of revocations to a quorum of multiple revocation issuers which may be different from the credential issuer.
@@ 2.1 Feature Comparison
We briefly round up on feature differences between the named selective disclosure cryptographic schemes, as shown in table 1, mainly distinguishing between four fundamental features:
@@
  • URG: Unregroupability
Where wip is mentioned, it means work in progress on adoption of bitstring status lists for unlinkable revocations.
The “unregroupability” feature refers to the fact that different presentations of different claims cannot be linked to each other (regrouped) as presented by the same holder, even when they are signed by the same issuer.
@@ 2.2 Key contributions
The cryptographic scheme described in this paper, named SD-BLS for brevity, implements all the properties of the SD-JWT scheme and proposes a novel cryptographic approach to similar data structures. Furthermore SD-BLS proposes novel anonymous cryptographic revocation flow for verifiable credentials, that aims to solve governance issues posed by status and revocation lists.
@@
Similarly to the SD-JWT and mDOC formats, SD-BLS produces an array of claims: the elements of the array are individually signed by the issuer. In SD-BLS the signature(s) replace the HMAC and still enable the holder to selectively disclose only certain signed credentials, and produce a proof of possession that minimizes private information given to verifiers.
@@
SD-BLS proposes a novel approach to credential revocation: the data published by the revocation issuer will produce cryptographic material that contains no information about the credential holders. The cryptographic revocation material allows anyone to verify if an SD-BLS proof produced by a credential holder has been revoked. The unlinkability, and thus anonymity of the cryptographic revocation, allows the revocation issuer to share revocations in public and allows anyone to verify if credentials have been revoked.
Multi-stakeholder governance of revocations
SD-BLS allows to introduce a new trusted party to the issuance phase: the revocation dealer. The dealer doesn’t need to know the content of the credential or the identity of an holder: its role is that of producing the revocation signature and distributing its secret key to a configurable range of revocation issuers. Later on a configurable quorum of issuers may reconstruct the revocation secret key to revoke a credential. This way the decision on a revocation is not delegated to a single peer, but to a multi-stakeholder governance that protects the holder from issuer corruption.
Fast revocation checks
SD-BLS shows good benchmark results on revocation checks: the main operation a verifier needs to do is an additional $\mathbb{G}_{2}$ multiplication. The cost of this operation grows linearly with the number of published revocations.
@@ 2.3 Applications
The focus of the EUDI-ARF specifications is identity documents: it defines mechanisms and data structures to issue a Personal Identification (PID) as, for instance, with digital driving licenses. Similarly, the US government is experimenting with W3C-VC and mDOC for cross-states interoperable driving licenses. The SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and anonymous revocation.
@@
  • A proof of possession can be published and be peer-verified on-chain in its private form without disclosing the credential contents.
  • A smart-contract may verify if one or more holder’s credentials match the requirement needed to process a transaction.
@@ 3.1 Notations and assumptions
We also require the notion of a cryptographic pairing. [11]
For the purpose of our protocol we will consider the Miller pairing $e:E_{T}\times E\to\mathbb{G}_{T}$, where $\mathbb{G}_{T}\subset\mathbb{F}_{p^{12}}$ is the subgroup containing the $n$-th roots of unity, and $n$ is the order of the groups $E$ and $E_{T}$.
For completeness we also recall the main properties of the map:
@@
    Bilinearity, i.e. given $P_{1},Q_{1}\in E_{T}$ and $P_{2},Q_{2}\in E$, we have
@@
    Non-degeneracy, meaning that for all $g_{1}\in E_{T},g_{2}\in E$, $e(g_{2},g_{1})\neq 1_{\mathbb{G}_{T}}$, the identity element of the group $\mathbb{G}_{T}$;
@@
    $E_{T}\neq E$, and moreover, that there exist no efficient homomorphism between $E_{T}$ and $E$.
@@ 3.2 Issuance
As for other well known algorithms BLS signing works following three main steps:
  • Key Generation phase.
    For an issuer who wants to sign a credential $m$, a secret key $sk$ is a random number chosen uniformly in $\mathbb{F}_{n}$, where $n$ is the order of the groups $\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T}$. The corresponding public key $pk$ is the element $sk\cdot G_{2}\in E_{T}$;
  • Signing phase.
    The credential $m$ is first hashed into the point $U\in E$; the related signature is then given by $\sigma=sk\cdot U$;
  • Verification phase.
    For an other user that wants to verify the authenticity and the integrity of the message $m$, it needs to
@@
The issuer create a credential as follow: given a claim $m$, it generate a new secret revocation key $rev$ with public key $r$. Let $\mathcal{H}$ be a cryptographic hash function, we compute:
$H=\mathcal{H}(m:r)$
then the issuer proceeds to sign with both its private key, and the revocation key generated above:
$\begin{split}r&=rev\cdot G_{2}\\ \sigma_{rev}&=sign(rev,H:r)\\ \sigma&=sign(sk,H:r)+\sigma_{rev}\\ \end{split}$
@@
$\mathcal{C}=\big{\{}\{H,r,\sigma,m\}:m\in claims\big{\}}$
At the end of this phase the holder is sent the signed claims to be stored in a private wallet, while the issuer stores a list of revocations as tuples formed by $\{H,rev\}$ into a private database that can be used later to issue revocations.
@@ 3.3 Presentation
Any presentation of a SD-BLS credential can simply omit the message $m$ to separate disclosure from verification as required by the specific context of a proof of possession and to satisfy privacy preserving design patterns for data minimization.
Basic Proof
A credential holder can choose any set of signed claims to present, and selectively disclose them into what we name “Basic Proof” presentation.
The holder can present a basic proof of possession $\{H,r,\sigma\}$ for each claim requested, or the complete set $\{H,r,\sigma,m\}$ for claims whose content must be disclosed.
However, Both forms of presentation of a basic proof are vulnerable to replay attacks: any verifier receiving such presentations can reuse them to impersonate the holder in any other session.
One Time Proof
To prevent replay attacks we exploit once again the homomorphic property of BLS signatures to add a sessions signature into the credential signature, then we add separate fields containing a signed timestamp and the public key of the session signature.
We name this presentation “One Time Proof”.
For each presentation the holder generates a key pair $(sk_{t},pk_{t})$ and a string $t$ containing session information, i.e. an expiration date or a pointer to the intended recipient. Then we compute:
$\begin{split}\sigma^{\prime}&=\sigma+sign(sk_{t},H:r)\\ \sigma_{t}&=sign(sk_{t},H:r:\sigma^{\prime}:t:pk_{t})\end{split}$
And obtain a presentation composed as follows:
$\{H,r,\sigma^{\prime},t,pk_{t},\sigma_{t}\}$
This schema is secure because it is impossible to reconstruct the holder information $\{H,r,\sigma\}$ necessary for a reply attack; indeed, given $\sigma^{\prime}$, an attacker that wants to retrieve $\sigma$ should be able to reconstruct the signature $\sigma_{t}$, but this is not possible without the knowledge of the secret key $sk_{t}$.
Furthermore the three attributes $t,\sigma_{t},pk_{t}$ cannot be modified or removed because $pk_{t}$ is necessary for the verification of signature $\sigma^{\prime}$. If an attacker tries to create a new valid presentation with freshly generated $sk_{t}^{\prime},pk_{t}^{\prime},t^{\prime}$ and $\sigma_{t}^{\prime}$, then the signature $\sigma^{\prime\prime}=\sigma^{\prime}+sign(sk_{t}^{\prime},H:r)$ can be verified using the public key $pk_{t}+pk_{t}^{\prime}$. However this public key does not verify the signature $\sigma_{t}^{\prime}$, and $\sigma_{t}^{\prime}$ can not be updated using the homomorphic property since it is the signature over the old $\sigma^{\prime}$. Thus any tampering of the presentation will lead to an invalid credential.
@@ 3.4 Verification
Credential verification is made by checking the presented issuer’s signature and revocation status of each claim. In order to verify the signature the credential issuer’s public key must be added to the revocation public key.
Basic proof
We consider the basic credential proof as a collection of tables of the following form:
$c=\{H,r,\sigma\}$
where $\sigma$ and $r$ are respectively the signature of the string $H$ and the revocation public key, optionally the claim value $m$ can be included.
We can check the validity of the presented claim computing the key:
@@
As proof of correctness consider that the signature $\sigma$ is given by:
$\begin{split}\sigma&=sign(A.sk,H:r)+sign(rev,H:r)\\ &=A.skU+revU\end{split}$
where $U$ is the mapping of the string $H$ in the group $G_{1}$.
@@
When the presentation contains the value $m$ it should also be checked that:
$H=\mathcal{H}(m:r).$
One Time Proof
If the verifier receives the credential locked to the current session (with replay attack protection) as $\{H,r,\sigma^{\prime},t,pk_{t},\sigma_{t}\}$ then it proceeds to
  • verifiy the information in $t$ (e.g. timestamp)
  • verify $\sigma_{t}$ with public key $pk_{t}$
  • verify $\sigma^{\prime}$ with the public key $A.pk+r+pk_{t}$
In case the verifier is presented with a one time disclosure that includes $m$ then it needs also to check its hashed value $H=\mathcal{H}(m:r)$.
This concludes the first verification phase. If the given presentation is valid, then the verifier should proceed to check the revocation status of the credential.
@@ 3.5 Revocation
Given an element $c=\{H,r,\sigma\}$ of the credential presentation, and a series of $rev$ elements found in any revocation list, we can verify a revocation by checking if the revocation public key presented by the holder matches:
@@
Note that, in the case the verifier received a credential in the “One Time Proof” form, the check for the revocation status does not change.
@@ 3.6 Threshold Revocation
Such a revocation dealer will be in contact with credential issuers for the signature of credentials: it will create the $rev$ revocation key while concealing it from them. The dealer then proceeds creating the $\sigma_{rev}=sign(rev,H:r)$ revocation signature and the $r=rev\cdot G_{2}$ public key, which will be communicated to the credential issuer to be aggregated into the signed credential (see section (3.2)).
The dealer will then proceed to split the secret revocation key into shares using a public verifiable secret sharing (PVSS [12]) implementation and distribute these shares to all revocation issuers. In order to issue a revocation, a configurable quorum of peers among the revocation issuers will need to reconstruct the secret revocation key and publish it.
@@ 4 Benchmarks
Figure 1: Speed comparison of issue and verify in SD-BLS
As a handy reference, we provide a benchmark of the same operations performed using the BBS+ in Zenroom on the same machine: results show that SD-BLS is one order of magnitude slower than BBS+, whose verification runs aproximately at the same speed as other operations.
Figure 2: Speed comparison of issue and verify in BBS+
Based on our benchmarks, the resulting data objects sizes in SD-BLS are:
  • Signed claim: 177 Bytes
  • Proof: 322 Bytes
@@
The computational cost of verifications grows linearly in presence of a cryptographic revocations list. We assume revocations are published as lists of secret revocation keys. Lab measurements of the time taken by a single proof verification process to operate on a growing number of revocations is shown in figure 3.
@@ 5.1 Security considerations
The revocations database is privacy and corruption sensitive (by our previous definition of issuer corruption) and it should be securely stored by each Issuer. This is mitigated by the adoption of threshold revocation. When using threshold revocation there is still the need for a revocation dealer who has no access to private informations, but may be dishonest and publish revocation keys collected during the process of credential issuance.
BLS signatures and the proof system obtained with credentials are considered secure by assuming the existence of random oracles [13], together with the decisional Diffie-Hellman Problem (DDH) [14], the external Diffie-Hellman Problem (XDH), and with the Lysyanskaya-Rivest-Sahai-Wol Problem (LRSW) [15], which are connected to the Discrete Logarithm. The future growth of quantum-computing technologies may be able to overcome the Discrete Logarithmic assumptions by qualitatively different computational means and SD-BLS may then be vulnerable to quantum-computing attacks. However this is speculative reasoning on what we can expect from the future.
@@
The $H$ component may be protected against brute-forcing attacks using hash collisions to forge a valid credential for a different message. A protection against this attack can be the adoption of a key-derivation function like Argon2 [18] on hash creation, which will add computational costs to issuance and verification.
Anyone with knowledge of $H,r$ can try to guess $m$ by appending known strings, which becomes trivial especially in case of boolean credential strings (i.e. “above18=true”). We recommend the credential issuer appends a nonce like $m:nonce$ before hashing and signing the message, which will be known by the holder and always disclosed when disclosing $m$.
@@ 5.2 Future development
EUDI-ARF dictates that the holder’s secret key generation and signatures must occur inside a trusted platform module (TPM). For mobile devices, this limits the secret keys and signatures to those offered, via proprietary APIs by the mobile OS, namely RSA (multiple flavours) and ECDSA on the secp256r1 curve. Currently the TPMs APIs supported by Android and iOS do not support BLS 12-381 key generation or signature.
@@
The SD-BLS math is fully compatible with ETH2.0 and can be computed inside an Ethereum VM. A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [22]) that want to authenticate peers using the selective disclosure of verifiable credentials instead of a single key based proof of possession.